OpenStudy (anonymous):

what is the completing the square procedure for factoring? I always get confused on these

OpenStudy (lgbasallote):

the process for completing the square: ax^2 + bx + c = 0 ax^2 + bx +(b/2)^2 = -c + (b/2)^2 (ax+b/2)^2 = -c + b^2/4 ax + b/2 = sqrt of (-c + b^2/4) ax = sqrt of(b^2/4 -c) - b/2 ax = sqrt of (b^2 -c)/2 - b/2 x = sqrt of (b^2 -c) -b all over 2a that's supposed to show the quadratic equation but it's still right